OpenStudy (anonymous):

Choose the point which is a solution to the system of inequalities: 4x + 3y > 12 x – y ≤ –2 (2, 1) (–4, 9) (–5, –6) (3, 7)

OpenStudy (across):

I will try the first point...\[4(2)+3(1)=8+3=11\not{>}12\]

OpenStudy (across):

... can you do the rest?

OpenStudy (anonymous):

Yeah Thanks, (: